OpenStudy (anonymous):

Rusty has a boat goes 12 mi. an hour. Find the current speed of the river if he can go 3 mi. upstream in the same time as he goes 5 mi. downstream.

OpenStudy (anonymous):

r t = d t = d/r Solve the following for c, the current speed.\[\frac{3}{12-c}\text{=}\frac{5}{12+c} \]c = 3 mph